OpenStudy (anonymous):

Find the points on the curve below at which the tangent is horizontal. Use n as an arbitrary integer. (Enter your answers as a comma-separated list.) y=cosx/(2+sinx)

OpenStudy (anonymous):

The tangent line is horizontal when its slope is 0, which is the same as when the derivative is 0. \[f(x)=\frac{\cos x}{2+\sin x}~~~\Rightarrow~~~f'(x)=\frac{-\sin x(2+\sin x)-\cos^2x}{(2+\sin x)^2}\] The derivative will be zero when its numerator is zero: \[\begin{align*}-\sin x(2+\sin x)-\cos^2x&=0\\ -2\sin x-\sin^2x-(1-\sin^2x)&=0\\ -2\sin x-1&=0\\ \sin x&=-\frac{1}{2}\\ x&=\frac{7\pi}{6}+2n\pi,~\frac{11\pi}{6}+2n\pi \end{align*}\]